# Lifshitz transitions and Weyl semimetals from a topological superconductor with supercurrent flow

### v1

@@ IV.3 Zero energy modes
Let us start from the Type-I Weyl phase, in the parameter regime (42). For $0<\mu<2w$, we find an arc connecting the projections of the nodes $\bm{k}_{W}^{0,\pm}$ and containing the point $Q=0$ and an arc connecting the projections of the nodes $\bm{k}_{W}^{\pi,\pm}$ containing $Q=\pi$, see Fig.3. Conversely, for $-2w<\mu<0$, the projection of the nodes $\bm{k}_{W}^{\pi,\pm}$ is closer to the point $Q=0$ than the projection of the nodes $\bm{k}_{W}^{0,\pm}$: in this situation, there is an arc connecting $\bm{k}_{W}^{0,\pm}$ across the BZ and an arc connecting $\bm{k}_{W}^{\pi,\pm}$ including the origin. In terms of the 1D model, the endpoints signal the transition between the topological and the trivial phase  [^1] Note that in the Weyl-I regime (42), if the denominator of $\lambda_{Q}$ is positive (topological phase), the argument of the square root in the numerator is also positive.. In Appendix E, we follow an established route for constructing Fermi arcs in Weyl Hamiltonian [66, 67, 68, 69]. Linearizing the Hamiltonian (33) in $k$, one finds localized eigenstates in the form
@@ IV.4 Comparison with the case of a $s$-wave superconductor
Let us now compare the Fermi surface closed circuits of the $p$-wave superconductor and $s$-wave superconductor. In the former case, $\mathcal{F}$ consists of the four curves that in Figs.4(a) delimit the electron and hole pockets. In particular, the ones highlighted as solid or dashed are given by
@@
respectively. Note that, at the crossing at the Weyl nodes, the solid and the dashed line are uniquely identified through the continuity of the eigenvectors related to the vanishing eigenvalues $E_{+}=E_{-}=0$. As a consequence, the Kitaev Fermi surface winds around the $k$-direction and therefore belongs to the $(1,0)$ topological class. In contrast, for the $s$-wave superconductor described, the BCS Fermi surface illustrated in Fig.5(a) is always localized and belongs to the $(0,0)$ class, which is homotopically equivalent to a point.
@@
Figure 5: The case of an $s$-wave superconductor. $(a)$ Fermi surface for the 2D BCS model where the superconducting phase modulation $Q$ is treated as a synthetic dimension. The Fermi surfaces surrounding the electron (cyan) and hole (pink) pockets do not extend along the entire Brillouin zone rather remains localized. The grey areas denoted the gapped regions and determined by the boundaries $Q^{\prime}$ and $\pi-Q^{\prime}$ while the remaining regions show the coexistence of Cooper pairs and unpaired electrons and holes. For BCS these pockets never touch each other as happens in the Kitaev model. As a consequence solid and dashed lines enclosing the electron and hole pockets are localized over a finite portion of the $k$-direction of the Brillouin zone, and belong to the homology class $(0,0)$. $(b)$ Upper (blue) and lower (red) bands of the BCS model, in this case the associated 2D fermionic model corresponds to a conventional semimetal.
@@ V $Q$-dependence of the current
The qualitatively different behavior of $I(Q)$ in the Type-I vs Type-II parameter ranges (45) and (46) is further highlighted in Fig.7, which displays the current (56) as a function of $Q$, at $\mu=0$, for different values of the superconducting order parameter $\Delta_{0}/w$. As one can see from panel $\left(a\right)$, for $\Delta_{0}>w$ the 1D Kitaev model is in the WSM Type-I parameter range and the current exhibits a smooth behavior as a function of $Q$, whereas for $\Delta_{0}<w$ the Kitaev chain enters the WSM Type-II phase and the cusps clearly appear in the current. Panel (b) shows a density plot of the current as a function of $(Q,\Delta_{0})$, where the horizontal dashed lines represents the four cuts shown in panel (a).
@@
which represents the current carrier by Cooper pairs with momentum $-2Q$ and charge $-2{\rm e}$. Similarly, for $\Delta_{0}\ll w$, the maximal current $I^{*}$ reached at the cusp is
@@
Effects of disorder. We have also analyzed the effects of disorder on the behavior of the current. Specifically, we have considered a disordered on-site potential, which we have accounted for by replacing the constant chemical potential in the Hamiltonian (1) with $\mu\rightarrow\mu_{j}=\mu+w\,r\,\rho_{j}$, where $r$ is the disorder strength parameter and $\rho_{j}\in\left[-1,1\right]$ is a uniformly distributed random number. Figure 8 shows the $Q$-dependence of the current for $\Delta_{0}=0.6$, $\mu=0$, for three values of disorder strength, $r=0$ (clean case), $r=0.5$ (moderate disorder) and $r=1$ (strong disorder). Each curve corresponds to an average over a large number of disorder realizations. As far as the disorder remains moderate, the cusp behavior of the current at $Q=Q^{*}$ (vertical dashed line) remains robust, and the only effect of disorder is to reduce a bit the magnitude of the current. Only for strong disorder the curve smoothens and the cusp disappears.
@@ VI Discussion and Conclusions
We have demonstrated that, by treating the $Q$-modulation as the wavevector related of an extra synthetic dimension, it is possible to establish a mapping between the stationary nonequilibrium state of the 1D superconductor and the ground state of a 2D WSM. The Lifshitz transitions emerging in the Fermi surface of the 2D model as a function of the model parameters identify different phases of the topological superconductor travelled by a current. Exploiting such a mapping, we have also constructed the expression (49) of the Fermi “arc” in the 2D WSM and the Majorana quasi-particle (50) in the corresponding current-biased 1D TS.
@@ Acknowledgements.
F.D. acknowledges financial support from the NEThEQS PRIN 2022 Project (code 2022B9P8LN) and fruitful discussions with Mario Trigiante. F.G.M.C. acknowledges financial support from ICSC—Centro Nazionale di Ricerca in High-Performance Computing, Big Data, and Quantum Computing (Spoke 7), funded by European Union—Next Generation EU.
@@ Appendix D Lattice model
$\mathcal{H}_{2D}$ $=$ $\sum_{m,n}\Bigg{\{}\frac{\Delta_{0}}{2}\Big{[}c_{A\left(n+1,m\right)}^{\dagger}c_{B\left(n,m\right)}+c_{B\left(n,m\right)}^{\dagger}c_{A\left(n+1,m\right)}$ (102)
$-c_{A\left(n,m\right)}^{\dagger}c_{B\left(n+1,m\right)}-c_{A\left(n+1,m\right)}^{\dagger}c_{B\left(n,m\right)}\Big{]}$
$+\frac{w}{2}\Big{[}c_{A\left(n,m+1\right)}^{\dagger}c_{A\left(n+1,m\right)}+c_{A\left(n+1,m\right)}^{\dagger}c_{A\left(n,m+1\right)}$
$-c_{B\left(n+1,m+1\right)}^{\dagger}c_{B\left(n,m\right)}-c_{B\left(n,m\right)}^{\dagger}c_{B\left(n+1,m+1\right)}\Big{]}$
$+\frac{\mu}{2}\left[c_{A\left(n,m\right)}^{\dagger}c_{A\left(n,m\right)}-c_{B\left(n,m\right)}^{\dagger}c_{B\left(n,m\right)}\right]\Bigg{\}}\;.$

### v2

@@ IV.3 Zero energy modes
Let us start from the Type-I Weyl phase, in the parameter regime (42). For $0<\mu<2w$, we find an arc connecting the projections of the nodes $\bm{k}_{W}^{0,\pm}$ and containing the point $Q=0$ and an arc connecting the projections of the nodes $\bm{k}_{W}^{\pi,\pm}$ containing $Q=\pi$, see Fig.3. Conversely, for $-2w<\mu<0$, the projection of the nodes $\bm{k}_{W}^{\pi,\pm}$ is closer to the point $Q=0$ than the projection of the nodes $\bm{k}_{W}^{0,\pm}$: in this situation, there is an arc connecting $\bm{k}_{W}^{0,\pm}$ across the BZ and an arc connecting $\bm{k}_{W}^{\pi,\pm}$ including the origin. In terms of the 1D model, the endpoints signal the transition between the topological and the trivial phase  [^1] Note that in the Weyl-I regime (42), if the denominator of $\lambda_{Q}$ is positive (topological phase), the argument of the square root in the numerator is also positive.. In Appendix E, we follow an established route for constructing Fermi arcs in Weyl Hamiltonians [67, 68, 69, 70]. Linearizing the Hamiltonian (33) in $k$, one finds localized eigenstates in the form
@@ IV.4 Comparison with the case of a $s$-wave superconductor
Let us now compare the Fermi surface closed circuits of the $p$-wave superconductor and $s$-wave superconductor. In the former case, $\mathcal{F}$ consists of the four curves that in Fig.4(a) delimit the electron and hole pockets. In particular, the ones highlighted as solid or dashed are given by
@@
respectively. Note that, at the crossing at the Weyl nodes, the solid and the dashed line are uniquely identified through the continuity of the eigenvectors related to the vanishing eigenvalues $E_{+}=E_{-}=0$. As a consequence, the Fermi surface in the Kitaev case winds around the $k$-direction and therefore belongs to the $(1,0)$ topological class. In contrast, for the $s$-wave superconductor described by the BCS model, the Fermi surface illustrated in Fig.5(a) is always localized and belongs to the $(0,0)$ class, which is homotopically equivalent to a point.
@@
Figure 5: The case of an $s$-wave superconductor. $(a)$ Fermi surface for the 2D BCS model where the superconducting phase modulation $Q$ is treated as a synthetic dimension. The Fermi surfaces surrounding the electron (cyan) and hole (pink) pockets do not extend along the entire Brillouin zone rather remains localized. The gray areas denote the gapped regions and determined by the boundaries $Q^{\prime}$ and $\pi-Q^{\prime}$ while the remaining regions show the coexistence of Cooper pairs and unpaired electrons and holes. These pockets never touch each other, differently from what happens in the Kitaev model. As a consequence, solid and dashed lines enclosing the electron and hole pockets are localized over a finite portion of the $k$-direction of the Brillouin zone, and belong to the homology class $(0,0)$. $(b)$ Upper (blue) and lower (red) bands of the BCS model, in this case the associated 2D fermionic model corresponds to a conventional semimetal.
@@ V $Q$-dependence of the current
The qualitatively different behavior of $I(Q)$ in the Type-I vs Type-II parameter ranges (45) and (46) is further highlighted in Fig.7, which displays the current (56) as a function of $Q$, at $\mu=0$, for different values of the superconducting order parameter $\Delta_{0}/w$. As one can see from panel $\left(a\right)$, for $\Delta_{0}>w$ the 1D Kitaev model is in the WSM Type-I parameter range and the current exhibits a smooth behavior as a function of $Q$, whereas for $\Delta_{0}<w$ the Kitaev chain enters the WSM Type-II phase and the cusps clearly appear in the current. Panel (b) shows a density plot of the current as a function of $(Q,\Delta_{0})$, where the horizontal dashed lines represent the four cuts shown in panel (a).
@@
which represents the current carried by Cooper pairs with momentum $-2Q$ and charge $-2{\rm e}$. Similarly, for $\Delta_{0}\ll w$, the maximal current $I^{*}$ reached at the cusp is
@@
Effects of disorder. We have also analyzed the effects of disorder on the behavior of the current. Specifically, we have considered a disordered on-site potential, which we have accounted for by replacing the constant chemical potential in the Hamiltonian (1) with $\mu\rightarrow\mu_{j}=\mu+w\,r\,\rho_{j}$, where $r$ is the disorder strength parameter and $\rho_{j}\in\left[-1,1\right]$ is a uniformly distributed random number. Figure 8 shows the $Q$-dependence of the current for $\Delta_{0}=0.6$, $\mu=0$, for three values of disorder strength, $r=0$ (clean case), $r=0.5$ (moderate disorder) and $r=1$ (strong disorder). Each curve corresponds to an average over a large number of disorder realizations. As far as the disorder remains moderate, the cusp behavior of the current at $Q=Q^{*}$ (vertical dashed line) remains robust, and the only effect of disorder is to reduce a bit the magnitude of the current. Only for strong disorder the curve becomes smooth and the cusp disappears.
@@ VI Discussion and Conclusions
We have demonstrated that, by treating the $Q$-modulation as the wavevector related of an extra synthetic dimension, it is possible to establish a mapping between the stationary nonequilibrium state of the 1D superconductor and the ground state of a 2D WSM. The Lifshitz transitions emerging in the Fermi surface of the 2D model as a function of the model parameters identify different phases of the topological superconductor crossed by a current. Exploiting such a mapping, we have also constructed the expression (49) of the Fermi “arc” in the 2D WSM and the Majorana quasi-particle (50) in the corresponding current-biased 1D TS.
@@ Acknowledgements.
F.D. acknowledges financial support from the NEThEQS PRIN 2022 Project (code 2022B9P8LN) funded by the Italian Ministry of University and Research and fruitful discussions with Mario Trigiante. F.G.M.C. acknowledges financial support from ICSC—Centro Nazionale di Ricerca in High-Performance Computing, Big Data, and Quantum Computing (Spoke 7), funded by European Union – Next Generation EU (Grant No. CN00000013).
@@ Appendix D Lattice model
$\mathcal{H}_{2D}$ $=$ $\sum_{m,n}\Bigg{\{}\frac{\Delta_{0}}{2}\Big{[}f_{A\left(n+1,m\right)}^{\dagger}f_{B\left(n,m\right)}+f_{B\left(n,m\right)}^{\dagger}f_{A\left(n+1,m\right)}$ (102)
$-f_{A\left(n,m\right)}^{\dagger}f_{B\left(n+1,m\right)}-f_{A\left(n+1,m\right)}^{\dagger}f_{B\left(n,m\right)}\Big{]}$
$+\frac{w}{2}\Big{[}f_{A\left(n,m+1\right)}^{\dagger}f_{A\left(n+1,m\right)}+f_{A\left(n+1,m\right)}^{\dagger}f_{A\left(n,m+1\right)}$
$-f_{B\left(n+1,m+1\right)}^{\dagger}f_{B\left(n,m\right)}-f_{B\left(n,m\right)}^{\dagger}f_{B\left(n+1,m+1\right)}\Big{]}$
$+\frac{\mu}{2}\left[f_{A\left(n,m\right)}^{\dagger}f_{A\left(n,m\right)}-f_{B\left(n,m\right)}^{\dagger}f_{B\left(n,m\right)}\right]\Bigg{\}}\;.$